# MACE-OFF: Short Range Transferable Machine Learning Force Fields for Organic Molecules

### v3

@@ Abstract
Classical empirical force fields have dominated biomolecular simulation for over 50 years. Although widely used in drug discovery, crystal structure prediction, and biomolecular dynamics, they generally lack the accuracy and transferability required for first-principles predictive modeling. In this paper, we introduce MACE-OFF, a series of short range transferable force fields for organic molecules created using state-of-the-art machine learning technology and first-principles reference data computed with a high level of quantum mechanical theory. MACE-OFF demonstrates the remarkable capabilities of short range models by accurately predicting a wide variety of gas and condensed phase properties of molecular systems. It produces accurate, easy-to-converge dihedral torsion scans of unseen molecules, as well as reliable descriptions of molecular crystals and liquids, including quantum nuclear effects. We further demonstrate the capabilities of MACE-OFF by determining free energy surfaces in explicit solvent, as well as the folding dynamics of peptides. Finally, we simulate a fully solvated small protein, observing accurate secondary structure and vibrational spectrum. These developments enable first-principles simulations of molecular systems for the broader chemistry community at high accuracy and relatively low computational cost.
@@ I Introduction
In this paper, we introduce a family of five purely local (i.e. short range), transferable bio-organic machine learning force fields which we refer to collectively as MACE-OFF. The force fields are parameterized for the ten most important chemical elements for organic chemistry: H, C, N, O, F, P, S, Cl, Br, and I. They are capable of accurately describing intra- and intermolecular interactions of neutral closed-shell systems. This enables the simulation of a wide range of chemical systems, from molecular liquids and crystals to drug-like molecules and biopolymers.
The models were validated on a number of tasks, including the prediction of small molecule torsion barriers, geometry optimizations, calculations of lattice parameters and enthalpies of formation of molecular crystals, and the calculation of the Raman spectra of molecular crystals, including nuclear quantum effects. We have also validated the models for predicting the densities and heats of vaporization of a range of molecular liquids. In particular, we study how well MACE-OFF reproduces the fundamental properties of water, including density and radial distribution functions. To further showcase the capabilities of the model, we computed the free energy surface and J-coupling parameters of alanine tripeptide (Ala3) in vacuum and explicit water, and simulated the folding of Ala15 at different temperatures. We also carried out an all-atom simulation of the protein crambin in explicit water (18 000 atoms), where we observed the correct secondary structure motifs and computed its vibrational spectrum, finding good agreement with other ML potentials and previously reported experimental data. Finally, we tested the computational speed of the current implementation in the LAMMPS and OpenMM simulation packages, demonstrating both strong and weak scaling.
While truly general purpose force fields for biomolecular modeling do need to include formally long-range Coulomb interactions, the purpose of this paper is to push the capabilities of short-range potentials as far as they can go, in part as preparation for including explicit polarizable electrostatic interactions in the future.
@@ II.1 The MACE architecture
$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
@@
where bold $\bm{l}\bm{m}$ denotes the $\nu$-tuple of $l$ and $m$ values and similarly to Equation (II.1), $\eta_{\nu}$ enumerates the number of possible couplings to create the features with equivariance $LM$. The maximum body-order is controlled by the parameter $\nu$ and is fixed at 3 (corresponding to 4-body terms, which includes the central atom) for all MACE-OFF models in this work.
@@ II.2 Training data
The core of the training set is version 1 of the SPICE dataset [51] with 95% of the data used for training and validation, and 5% set aside for testing. Test/train splitting was performed at the molecule level, ensuring conformers of the same molecule do not appear in both train/validation and test sets. Table 1 summarizes the training and test sets. The MACE-OFF models are trained to reproduce the energies and forces computed at the ωB97M-D3(BJ)/def2-TZVPPD level of quantum mechanics [52, 53, 54, 55, 56], as implemented in the PSI4 software package [57]. We have used a subset of SPICE that contains the ten chemical elements H, C, N, O, F, P, S, Cl, Br, and I, and has a neutral formal charge (see Table 2 in Ref. [51] for element coverage in the dataset). We have also removed the ion pairs subset. Overall, we used about 85% of the full SPICE dataset. The geometries in the SPICE dataset have been generated by running molecular dynamics simulations using classical force fields [58], performed at both 300K and 500K, and sampling maximally different conformations from the resulting trajectories [51].
Table 1: Summary of training and test sets. The columns PubChem to Solvated Amino Acids correspond to the SPICE dataset [51].
@@
After fitting the small MACE model (see Section II.3) and observing the training errors, we noticed the presence of outliers in the dataset. This is probably caused by errors in the underlying electronic structure calculations, some of which have been documented on the SPICE GitHub repository.***https://github.com/openmm/spice-dataset To confirm this, we also re-evaluated a selection of the configurations that had the highest error, using the level of DFT used in SPICE, and found that for about a third of the configurations, the recomputed energies and forces agreed well with the MACE prediction and not with the original DFT labels. To purify the dataset, we removed from the training set the configurations that had a maximum force error greater than 2 eV/Å. This meant the removal of just 808 configurations, many of which contained heavy elements, in particular phosphorus and iodine, that might have a more challenging electronic structure.
In order to understand the role additional data plays in the performance of the model, we also investigated the difference in performance due to inclusion of additional configurations released as part of the second version of SPICE [63]. The updated dataset, which we incorporate into the MACE-OFF24(M) model (see Section II.3), contains among others, an additional 208,000 configurations comprising solvated PubChem molecules and amino acid-ligand pairs. The former were sampled from molecular dynamics using the same protocol as in the original dataset, whilst the amino acid-ligand pairs were extracted from structures in the PDB. This release also contained 1000 water clusters, containing up to 90 atoms. These were also included and used to complement the set of water clusters we had already included in the extended dataset.
In common with the first version of SPICE, close inspection of this new set revealed several nonphysical geometries with correspondingly high DFT forces. These configurations were removed by applying a maximum force filter of 15 eV/Å. We also observed that many phosphorous-containing compounds were present among the outliers, as well as configurations with highly distorted geometries, which also suggest challenging electronic structure. We also found that many configurations had forces that do not sum to zero. To avoid the possibility of these configurations poisoning the dataset, we applied a total force filter of 0.1 eV/Å. In total this removed 10,372 configurations from the SPICE v2 subsets.
@@ II.3 Model details
The MACE model has parameters that enable systematic control of model expressivity (and subsequent accuracy) against computational cost. In this paper, we present three variants of the MACE-OFF23 model, a small, a medium, and a large one, denoted in the text as MACE-OFF23(S), MACE-OFF23(M) and MACE-OFF23(L) respectively. The models get more accurate with size, but also have an increasing computational cost. The small model is well suited for large-scale simulations, the medium model offers a good balance of speed and accuracy, and the large model is best used for small systems or when the highest possible accuracy is desirable.
@@
23(S) 23(M) 23(M)b 23(L) 24(M)
Cutoff radius (Å) 4.5 5.0 6.0 5.0 6.0
Chemical channels 96 128 128 192 128
@@
max L (see Eq. (5)) 0 1 1 2 1
SPICE version 1 1 1 1 2
The hyperparameters of the models are displayed in Table 2. The MACE-OFF23(S) model has 96 channels, $L=0$ invariant messages, and a cutoff of 4.5 Å. The MACE-OFF23(M) model has 128 channels and $L=1$ messages, and the MACE-OFF23(L) has 192 channels and $L=2$ equivariant messages. Medium and large models have a cutoff of 5.0 Å in each layer. In light of the additional configurations added to the SPICE dataset during this work, we present two additional models that improve upon the performance of the original three MACE-OFF23 models. MACE-OFF23(M)b is identical to the medium model except for the increased 6 Å cutoff, which we find is crucial for recovering certain properties of condensed phase systems. MACE-OFF24(M) retains the extended 6 Å cutoff and includes additional configurations from SPICE v2 dataset [63]. MACE-OFF23(M)b can therefore be viewed as an ablation of the MACE-OFF24(M) model, enabling insight into which properties are driven by the extended cutoff and those that additionally require the additional nonbonded interaction datasets. All models were trained using the PyTorch [64] implementation of MACE, available on https://github.com/ACEsuit/mace. For more information about the training, please refer to the SI.
II.4 MACE-OFF23-$\mu$
In addition to the series of MACE-OFF force field models, we also present MACE-OFF23-$\mu$, which predicts the total dipole moment of molecules and can be used to estimate vibrational spectra based on the autocorrelation of the dipole moment. The model architecture is identical to the MACE architecture described above, with the only difference being the readout function, which produces an $L=1$ tensor for each atom instead of the atomic site energies. Given the equivariant order of the readout, the smallest possible model corresponds to the ‘medium’ architecture, which is the one considered in this work. Other than the readout’s equivariant order, all hyperparameters match those of the MACE-OFF23(M) model. MACE-OFF23-$\mu$ was trained on the subset of SPICE v1 for which a total dipole moment was available, resulting in the exclusion of the QMugs and water subsets.
@@ III.1 Extended SPICE Test Set
First, we look at the pointwise errors of the energy and force predictions on a held-out test set for each of the three baseline MACE-OFF23 models. Figure 1 shows the per-atom energy, force and intermolecular force root-mean-square errors (full statistics are provided in SI Table 4). As the size of the model increases, the models gradually become more accurate, with the large model generally achieving errors of around 0.5–1.0 meV/atom and 15–20 meV/Å, well below the 1 kcal/mol (43 meV) chemical accuracy limit for the drug-like organic molecules studied here.
Figure 1: Test set root mean square errors (RMSE). Errors in the MACE-OFF23 models compared to the underlying DFT reference data, highlighting the relative accuracy of the three models. Bottom panels show specifically inter-molecular force errors compared to overall DFT inter-molecular force magnitudes (RMS) and note the absence of results for subsets that are composed of single molecules (i.e. DES370K Monomers, Dipeptides, QMugs, Tripeptides). Note that for subsets comprising only single molecule configurations (i.e. DES370K Monomers, Dipeptides, QMugs, Tripeptides) inter-molecular contributions are expected to be zero. The slight deviation from zero arises because DFT forces do not obey translational and rotational symmetries with sufficient accuracy, while MACE models do (check SI section B for details).
@@
The bottom panel of Figure 1 shows specifically the intermolecular force errors, which were obtained by separating the force contributions to molecular translations and rotations (see SI and Ref. [65] for details). These interactions are crucial as they underpin the thermodynamics and transport properties in the organic condensed phase. MACE-OFF23(M) and MACE-OFF23(L) both yield similar errors of around 5–15 meV/Å, which are about 1.5–3 times smaller than total force errors, and about 5–10% of the typical DFT intermolecular force magnitude (RMS). For context, relative intramolecular force errors, which are roughly equivalent to total force errors, are around 1–2%. The latter are routinely used across the literature to benchmark ML models, however, we advocate that intermolecular RMSE errors are a more appropriate metric to assess model accuracy for organic condensed phase applications. The results here demonstrate the remarkable accuracy of MACE models in predicting these subtle but crucial interactions. We will investigate how these energy and force errors translate into condensed phase physical property predictions in later Sections.
Figure 2: Test set prediction and root mean square errors (RMSE). Violin plots of the reference (blue) and predicted (red) distributions of the dipole moment predicted using MACE-OFF23-$\mu$. The RMSEs are reported above the violin plots for the individual test sets.
We also assess the accuracy of the MACE-OFF23-$\mu$ model on the test sets, evaluating its effectiveness for predicting dipole moments and infrared spectra of organic molecules. Figure 2 summarizes this performance by examining the alignment between the violin distributions of the reference and predicted total dipole moment values, as well as the RMSEs. The small RMSEs and strong alignment of the distributions indicate high fidelity across the test datasets for which reference dipole moment values were available. These results suggest promising performance for zero-shot predictions of the dipole moments (and infrared spectra) of unseen organic molecules, as demonstrated in a recent benchmark on out-of-distribution drug-like molecules from the NIST database [66].
@@ III.2.1 TorsionNet-500
The top panel of Figure 3 summarizes the results for the TorsionNet-500 dataset [70]. This dataset contains torsion drives of 500 different molecules, selected to cover a wide range of pharmaceutically relevant chemical space. The original data set was reported at the B3LYP/6-31G(d) level of DFT theory. For consistency with the SPICE dataset, we recreated the torsion profiles using the DFT setting of SPICE, which uses a higher level of theory and basis set (SI Section C).
The first panel shows an example of a torsion drive, indicating the complex energy profile that the MACE models are able to capture closely, including geometries far from equilibrium. The center panel shows the mean barrier height error of a number of representative models, comparing the Sage classical empirical force field [21], a semi-empirical quantum mechanical method GFN2-xTB [23], a recent transferable machine learning force field AIMNet2 [8], and the three MACE-OFF23 models. Again, systematic improvements in accuracy with MACE model size are observed, with medium and large models, in particular, achieving errors of around 0.25 kcal/mol compared to the reference method. The AIMNet2 model achieves comparable accuracy to the small MACE-OFF23 model. A similar conclusion can be drawn from the comparison of the molecular geometries by looking at the root mean squared deviation of the atomic positions averaged over the full torsion scans, as indicated by the top right panel of Figure 3. It shows that MACE-optimized geometries have a deviation of about 0.025 Å, meaning they are almost indistinguishable from DFT-optimized structures. It is important to note that the other models were trained to different levels of DFT, which might also contribute to the observed differences.
@@ III.2.3 Infrared spectroscopy of a drug molecule
To assess the accuracy of the MACE-OFF23 and MACE-OFF23-$\mu$ models for finite temperature properties of molecules, we study the infrared spectrum of an isolated paracetamol molecule at room temperature. This is a stringent test, as reproducing the line positions and their relative intensities requires an accurate treatment of both intramolecular interactions and the dipole moment in out-of-equilibrium configurations. As shown in Figure 4, we estimate the classical and approximate quantum IR spectra [71] using the MACE-OFF23(M) and MACE-OFF23-$\mu$ models, and compare them directly with experimental data at 293 K [72]. For low and intermediate frequencies (up to around 2000 cm-1), the classical and quantum spectra agree with each other and with the experimental data, allowing identification of all collective modes. For the high-frequency O–H and N–H modes, we observe that the classical spectra are shifted to the blue due to the absence of anharmonic zero-point fluctuations, which typically soften these bonds. A quantum treatment of nuclear dynamics results in good agreement with the experimental data, with the exception of a net blue shift, likely due to the limitations of the reference density functional. We also note a slight discrepancy in the relative intensities of the high-frequency stretching with respect to the lower-frequency modes. This could result from a limitation of the reference density functional or the MACE-OFF23-$\mu$ model, which we aim to investigate in a future study.
@@
Figure 4: Infrared (IR) spectrum of paracetamol molecule. The classical (blue dashed) and quantum (blue solid) IR spectrum at 293 K using the MACE-OFF23(M) and MACE-OFF23-$\mu$ models by propagating the system and estimating the time correlation function of the time derivative of the total dipole moment. Quantum nuclear effects are incorporated using path-integral coarse-grained simulations (PIGS) [71]. The experimental data are taken from Ref. 72.
III.3 Molecular crystals
In the following, we demonstrate the ability of the MACE-OFF force fields to simulate the vibrational and thermal properties of molecular crystals.
III.3.1 Vibrational spectroscopy of paracetamol
Raman spectroscopy is one of the most widely used techniques for characterizing molecular crystals. Unlike IR spectroscopy, which only detects vibrational modes that distort dipoles, Raman spectroscopy is more sensitive to collective modes governed by weak, non-bonded interactions in a broad range of molecular materials. The low-frequency region of the Raman spectrum (e.g., the THz regime) gives a vibrational fingerprint of the intermolecular interactions. Thus, it is widely used to differentiate between polymorphs of molecular crystals. Meanwhile, the high-frequency Raman spectrum probes intramolecular modes and their coupling to low-frequency modes.
Here, we test the ability of the MACE-OFF23(S) model to predict the Raman spectrum of the “Form II” polymorph of paracetamol. To compare MACE-OFF23(S) directly with experiments, we rigorously incorporate quantum nuclear and non-Condon effects using our ML-aided framework [73]. We incorporate quantum nuclear effects by fitting an effective potential energy surface [71] (using MACE) to calculate quantum nuclear corrections to the MACE-OFF23(S) model within the path-integral coarse-grained simulations (PIGS) method. To incorporate non-Condon effects, we fit a separate equivariant MACE model to the first-principles polarizability of paracetamol polymorphs (data taken from Refs. 74, 75). The remaining steps of our quantum nuclear simulations mimic an entirely classical calculation involving a $NVE$ molecular dynamics simulation, prediction of the isotropic and anisotropic components of the polarizability tensor, and calculation of their time correlation functions. For further details we refer the reader to Ref. 73.
As shown in Figure 5, we predict both the high- and low-frequency regions of the Raman spectrum of paracetamol form II with an overall good agreement with the experimental band positions [76]. Since the experiment captures the Raman spectra along different crystal directions while we estimate the “powder” Raman spectrum [75], we do not compare band intensities. We find that the classical predictions (based on MACE classical MD) are consistently shifted with respect to the experiment. At the same time, quantum nuclear predictions (encoded by the PIGS method) play an essential role in improving the agreement between theory and experiments. Moreover, we note that a broad band at around 3300 $\text{cm}^{-1}$ is only captured at the level incorporating quantum nuclear effects. The low-frequency modes also agree semi-quantitatively with the experiments and do not require a quantum nuclear description. The 2–3 % overall shift in the vibrational frequencies notwithstanding, the MACE-OFF23(S) combined with quantum nuclear effects and non-Condon effects shows promising capabilities for the characterization of molecular crystals.
Figure 5: Powder Raman spectrum of paracetamol form II [77]. Spectrum computed at ambient conditions using the MACE-OFF23(S) model for the potential energy surface, a MACE model of the polarizability, and a MACE model that incorporates quantum nuclear effects on the potential energy surface using the PIGS approach [71]. The black lines represent experimentally determined band positions [76], scaled by 3% to aid visual comparison with the predicted spectra.
III.3.2 Lattice enthalphies

### v4

@@ Abstract
Classical empirical force fields have dominated biomolecular simulation for over 50 years. Although widely used in drug discovery, crystal structure prediction, and biomolecular dynamics, they generally lack the accuracy and transferability required for first-principles predictive modeling. In this paper, we introduce MACE-OFF, a series of short range transferable force fields for organic molecules created using state-of-the-art machine learning technology and first-principles reference data computed with a high level of quantum mechanical theory. MACE-OFF demonstrates the remarkable capabilities of short range models by accurately predicting a wide variety of gas and condensed phase properties of molecular systems. It produces accurate, easy-to-converge dihedral torsion scans of unseen molecules, as well as reliable descriptions of molecular crystals and liquids, including quantum nuclear effects. We further demonstrate the capabilities of MACE-OFF by determining free energy surfaces in explicit solvent, as well as the folding dynamics of peptides and nanosecond simulation of a fully solvated protein. These developments enable first-principles simulations of molecular systems for the broader chemistry community at high accuracy and relatively low computational cost.
@@ I Introduction
In this paper, we introduce a family of purely local (i.e. short range), transferable bio-organic machine learning force fields which we refer to collectively as MACE-OFF. The force fields are parameterized for the ten most important chemical elements for organic chemistry: H, C, N, O, F, P, S, Cl, Br, and I. They are capable of accurately describing intra- and intermolecular interactions of neutral closed-shell systems. This enables the simulation of a wide range of chemical systems, from molecular liquids and crystals to drug-like molecules and biopolymers.
The models were validated on a number of tasks, including the prediction of small molecule torsion barriers, geometry optimizations, calculations of lattice parameters and enthalpies of formation of molecular crystals, and the calculation of the Raman spectra of molecular crystals, including nuclear quantum effects. We have also validated the models for predicting the densities and heats of vaporization of a range of molecular liquids. In particular, we study how well MACE-OFF reproduces the fundamental properties of water, including density and radial distribution functions. To further showcase the capabilities of the model, we computed the free energy surface and J-coupling parameters of alanine tripeptide (Ala3) in vacuum and explicit water, simulated the folding of Ala15 at different temperatures, and carried out an all-atom simulation of the protein crambin in explicit water (18 000 atoms). Finally, we tested the computational speed of the current implementation in the LAMMPS and OpenMM simulation packages, demonstrating both strong and weak scaling.
While truly general purpose force fields for biomolecular modeling require explicit treatment of long-range Coulomb interactions, the purpose of this paper is to push the capabilities of short-range potentials as far as they can go, in part as preparation for inclusion of explicit polarizable electrostatic interactions in the future.
@@ II.1 The MACE architecture
×$\qquad\qquad\times Y^{m_{1}}_{l_{1}}(\bm{\hat{r}}_{ij})h^{(t)}_{j,kl_{2}m_{2}}$ (2)
@@
where bold $\bm{l}\bm{m}$ denotes the $\nu$-tuple of $l$ and $m$ values and similarly to Eq. (II.1), $\eta_{\nu}$ enumerates the number of possible couplings to create the features with equivariance $LM$. The maximum body-order is controlled by the parameter $\nu$ and is fixed at 3 (corresponding to 4-body terms, which includes the central atom) for all MACE-OFF models in this work.
@@ II.2 Training data
[…]
Table 1: Summary of training and test sets. The columns ‘PubChem’ to ‘Solvated Amino Acids’ correspond to the original SPICE dataset [51].
@@
In order to understand the role additional data plays in the performance of the model, we also investigated the difference in performance due to inclusion of additional configurations released as part of the second version of SPICE [63]. The updated dataset, which we incorporate into the MACE-OFF24(M) model (see Section II.3), contains among others, an additional 208,000 configurations comprising solvated PubChem molecules and amino acid-ligand pairs. The former were sampled from molecular dynamics using the same protocol as in the original dataset, whilst the amino acid-ligand pairs were extracted from structures in the PDB. This release also included 1000 water clusters, containing up to 90 atoms. These were also added and used to complement the set of water clusters already included in the extended dataset.
@@ II.3 Model details
The MACE model has parameters that enable systematic control of model expressivity (and subsequent accuracy) against computational cost. In this paper, we present three variants of the MACE-OFF23 model, a small, a medium, and a large one, denoted in the text as MACE-OFF23(S), MACE-OFF23(M) and MACE-OFF23(L) respectively. The models get more accurate with size, but also have an increasing computational cost. The small model is well suited for large-scale simulations, the medium model offers a good balance of speed and accuracy, and the large model is best used for small systems or when the highest possible accuracy is desirable. MACE-OFF24(M) is an update to the medium model that uses an extended 6 Å cutoff and includes additional configurations from version 2 of the SPICE dataset [63]. The hyperparameters of the models are displayed in Table 2. All models were trained using the PyTorch [64] implementation of MACE, available at https://github.com/ACEsuit/mace. More information about the training is provided in Section S1.
@@
23(S) 23(M) 23(L) 24(M)
Cutoff radius (Å) 4.5 5.0 5.0 6.0
Chemical channels 96 128 192 128
@@
max L (see Eq. (5)) 0 1 2 1
SPICE version 1 1 1 2
@@ III.1 Extended SPICE Test Set
First, we look at the pointwise errors of the energy and force predictions on a held-out test set for each of the three baseline MACE-OFF23 models. Figure 1 shows the per-atom energy, force and intermolecular force root-mean-square errors (full statistics are provided in Section S2). As the size of the model increases, the models gradually become more accurate, with the large model generally achieving errors of around 0.5–1.0 meV/atom and 15–20 meV/Å, well below the 1 kcal/mol (43 meV) chemical accuracy limit for the drug-like organic molecules studied here.
Figure 1: Test set root mean square errors (RMSE). Errors in the MACE-OFF23 models compared to the underlying DFT reference data, highlighting the relative accuracy of the three models. Bottom panels show specifically inter-molecular force errors compared to overall DFT inter-molecular force magnitudes (RMS). Note that for subsets comprising only single molecule configurations (DES370K Monomers, Dipeptides, QMugs, Tripeptides) inter-molecular contributions are expected to be zero. The slight deviation from zero arises because DFT forces do not obey translational and rotational symmetries with sufficient accuracy, while MACE models do.
@@
The bottom panel of Figure 1 shows specifically the intermolecular force errors, which were obtained by separating the force contributions to molecular translations and rotations (see Section S2 and Ref. [65] for details). These interactions are crucial as they underpin the thermodynamics and transport properties in the organic condensed phase. MACE-OFF23(M) and MACE-OFF23(L) both yield similar errors of around 5–15 meV/Å, which are about 1.5–3 times smaller than total force errors, and about 5–10% of the typical DFT intermolecular force magnitude (RMS). For context, relative intramolecular force errors, which are roughly equivalent to total force errors, are around 1–2%. The latter are routinely used across the literature to benchmark ML models, however, we advocate that intermolecular RMSE errors are a more appropriate metric to assess model accuracy for organic condensed phase applications. The results here demonstrate the remarkable accuracy of MACE models in predicting these subtle but crucial interactions. We will investigate how these energy and force errors translate into condensed phase physical property predictions in later Sections.
@@ III.2.1 TorsionNet-500
The top panel of Figure 2 summarizes the results for the TorsionNet-500 dataset [69]. This dataset contains torsion drives of 500 different molecules, selected to cover a wide range of pharmaceutically relevant chemical space. The original dataset was reported at the B3LYP/6-31G(d) level of DFT theory. For consistency with the SPICE dataset, we recreated the torsion profiles using the DFT setting of SPICE, which uses a higher level of theory and basis set (Section S3).
The first panel shows an example of a torsion drive, indicating the complex energy profile that the MACE models are able to capture closely, including geometries far from equilibrium. The center panel shows the mean barrier height error of a number of representative models, comparing the Sage classical empirical force field [21], a semi-empirical quantum mechanical method GFN2-xTB [23], a recent transferable machine learning force field AIMNet2 [8], and the three MACE-OFF23 models. Again, systematic improvements in accuracy with MACE model size are observed, with medium and large models, in particular, achieving errors of around 0.25 kcal/mol compared to the reference method. The AIMNet2 model achieves comparable accuracy to the small MACE-OFF23 model. A similar conclusion can be drawn from the comparison of the molecular geometries by looking at the root mean squared deviation of the atomic positions averaged over the full torsion scans, as indicated by the top right panel of Figure 2. It shows that MACE-optimized geometries have a deviation of about 0.025 Å, meaning they are almost indistinguishable from DFT-optimized structures. It is important to note that the other models were trained to different levels of DFT, which might also contribute to the observed differences [70].
@@ III.2.3 Infrared spectroscopy of a drug molecule
To assess the accuracy of the MACE-OFF23(M) model for finite temperature properties of molecules, we study the infrared spectrum of an isolated paracetamol molecule at room temperature. Molecular dipole moments are predicted using the MACE architecture, trained on the SPICE dataset, and quantum nucelear effects are incorporated using path-integral coarse-grained simulations (PIGS) [71]. Full details are provided in Section S8. This is a stringent test, as reproducing the line positions and their relative intensities requires an accurate treatment of both intramolecular interactions and the dipole moment in out-of-equilibrium configurations. As shown in Figure 3, we estimate the classical and approximate quantum IR spectra [71], and compare them directly with experimental data at 293 K [72]. For low and intermediate frequencies (up to around 2000 cm-1), the classical and quantum spectra agree with each other and with the experimental data, allowing identification of all collective modes. For the high-frequency O–H and N–H modes, we observe that the classical spectra are blue-shifted due to the absence of anharmonic zero-point fluctuations, which typically soften these bonds. A quantum treatment of nuclear dynamics results in good agreement with the experimental data, with the exception of a net blue shift and a slight discrepancy in the relative intensities of the high-frequency stretching with respect to the lower-frequency modes, likely due to the limitations of the reference density functional.
@@
Figure 3: Infrared (IR) spectrum of a paracetamol molecule. The classical (blue dashed) and quantum (blue solid) IR spectrum at 293 K using the MACE-OFF23(M) model is computed by propagating the system and estimating the time correlation function of the time derivative of the total dipole moment. The experimental data are taken from Ref. 72.
III.3 Molecular crystals and organic liquids
In the following, we demonstrate the ability of the MACE-OFF force fields to simulate the vibrational and thermal properties of molecular crystals. Based on relative accuracy and computational expense, in what follows we present data only for the medium MACE-OFF models. Additional experiments, where either the higher throughput of the small model or the additional accuracy of the large model are desired, are reported in the Supporting Information.
III.3.1 Lattice enthalphies